# Particles before symmetry

### v4

@@ 5.3 Composite fibres and the Standard Model
From the PFB-POV this is no defect: it is essential to the standard strategy of assigning each particle type to a section of some associated bundle, without attempting to reconstruct the full gauge group from the geometry of any single fibre. Since it has no mandate to recover symmetry from geometry, the PFB-POV naturally allows a loose fit, or slack, between the structure group $G$ of $P$ and the geometry of the (composite) associated vector bundles that represent the particles. This slack poses no internal difficulty, for the PFB-POV treats the two ingredients—the symmetry and the vector bundle structure—merely as jointly posited and mutually consistent. [^31] If you think that is not how we actually use the principal-associated bundle formalism, I agree; see Section 5.4 below.
But if one nonetheless tries to recover the group from the automorphisms of the associated bundles corresponding to each particle’s label, not only would one fail to get the right verdict; in general, the recovery would not even be consistent.
@@
For concreteness, take $V\simeq{\mathbb{C}}^{3}\otimes{\mathbb{C}}^{2}$, $G=SU(3)\times SU(2)$ and consider two representations: $\rho_{1}=\mathbf{3}\otimes\mathbf{1}$ (a colour triplet, singlet under $SU(2)$), and $\rho_{2}=\mathbf{1}\otimes\mathbf{2}$ (a weak doublet, singlet under $SU(3)$). Thus in this case we would recover $G^{\prime}_{1}=SU(3)$ from one bundle, and $G^{\prime}_{2}=SU(2)$ from the other. What we would like, of course, is $SU(3)\times SU(2)$; but the product structure is nowhere to be found at the level of any single associated bundle. Naively, taking the product of automorphisms groups of the associated vector bundles won’t work either: by merely adding another particle whose representation labels included, say, $\mathbf{2}$ for $SU(2)$, we would get repetition, e.g. $SU(2)\times SU(2)$.
The VB-POV avoids this difficulty by shifting the level at which geometry fixes symmetry. Each gauge-group factor must arise as the automorphism group of a fundamental (or atomic) vector bundle, not of the composite associated bundles corresponding to particle multiplets with varying representation labels, or quantum numbers. These fundamental bundles serve as the basic building blocks: from them, one can reconstruct each factor ($SU(3)$, $SU(2)$, etc.) and then deduce their actions on the tensor products and direct sums that, in the VB-POV, characterise the various particles. [^32] Of course, particles may have no component in a given fundamental direction. At that composite stage, however, the group is no longer straightforwardly recoverable from geometry.
Fortuitously, in our world the symmetry group realised in Nature happens to fit the VB-POV’s stricter account of symmetry quite snugly. It need not have done so.
5.4 Methodological closure: why tight identification is a virtue
To close this Section, let me underscore this last point with a simple case. Above I gave several examples where the VB-POV would fail. The simplest example is a world in which the relation between charges of some type is irrational.
[…]
As a physical theory, a non-trivial connection $\varpi$ carries its own physical observables, independently of the behavior of matter. Thus one could, in principle, have independent evidence for a symmetry group inferred from $\varpi$, alongside a geometrical structure for the fermions invariant under a larger automorphism group, also independently physically significant. In such a world there would be more structure preserved under parallel transport—for instance an orientation—than the particles intrinsically possess (see footnote 33.)
But that’s not our world; at least it is not our Standard Model.
[…]
The slack-world scenario is implausible; but its very implausibility points to a tacit assumption already built into our use of principal–associated formalisms, which in practice never exploit the full flexibility of the PFB-POV. The assumption is precisely that such formalisms ultimately rest on theunderlying fundamental (or atomic) vector bundles, from which the total gauge group is written as:
$G=\prod G_{i},\quad G_{i}\simeq\rho(G_{i})\simeq\mathsf{Aut}(V_{i}).$ (5.15)
In other words, the standard use of principal and associated fibre bundles often tacitly presupposes the commitments of the geometry-first formulation of gauge theory.
This tacit assumption likely stems from an implicit empirical argument: that we cannot, even within the PFB-POV, obtain independent evidence for the connection and for the matter fields. But I resist such an argument, along with the tacit assumption. In the PFB-POV, $\mathsf{Aut}(V)$ could, in principle, be revealed by methods that do not rely on the principal connection or on parallel transport—for example, through the structure of bound states or potential terms. But I need not get bogged down in empirical or epistemological questions about what kind of experiment could in principle do this, or what would count as independent evidence for a section of $E$ and a connection on $P$. The point is simply that the gauge-invariant content of the principal connection is part of the physical apparatus of a gauge theory formulated in the PFB-POV. In that framework, the relevant group has physical significance and can differ from the automorphism group of the matter fields.
Moreover, physicists are not always consistent about this tacit assumption—which is why it is so hard to tease it out from the literature. Textbooks often begin with pedagogical, simplified cases where there is only one compact, semi-simple group and one vector space admitting the fundamental representation, as in the VB-POV. Indeed, this is how I introduced principal and associated vector bundles in Section LABEL:sef:PFB.
Yet I have never encountered in the literature an explicit mandate for always constructing principal and associated bundles “piecewise” in the manner the VB-POV advocates. If such a mandate exists, it remains tacit, and the literature is inconsistent about whether it is even intended. For after describing the relationship between $P$ and $E$ in the simple case where $P=L(E)$, where $G\simeq\rho(G)\simeq\mathsf{Aut}(V)$, expositions usually move straight to composite cases, where particles are assigned representation labels defining associated vector bundles under a product group such as $SU(3)\times SU(2)\times U(1)$ in the Standard Model.
For instance, \@BBOPcitep\@BAP\@BBN(Skinner, 2007)\@BBCP is among the textbooks that most clearly illustrate this line of exposition. Here we find perhaps the closest thing to an articulation of the tacit assumption—but he then goes on to deny it. On page 141 he writes: “Vector bundles are of relevance to physics because a charged matter field is a section of an associated vector bundle.” So far, his associated vector bundles can carry different quantum numbers, and so are not fundamental, in my sense; for that they would only be allowed to carry the fundamental representation, when it exists and is unique. But after constructing principal fibre bundles with matrix Lie groups from frame bundles of a vector space—i.e. assuming the fundamental representation and the validity of (5.3)—he says (ibid., p. 142): “The most common Lie groups that arise in physics are indeed matrix Lie groups [of that sort], so the two viewpoints [PFB and VB] are equivalent.” Here we see a restriction of the comparison to the cases that commonly arise in physics. I would resist saying they are ‘equivalent’: I take him to mean that the VB-POV can account for the cases of interest in physics. However, regarding which picture is more fundamental, he adds (ibid., p. 142): “However, in some exotic theories (especially string theory and some grand unified theories) exceptional Lie groups such as $E_{6}$ play an important role, so the fundamental picture is really that of principal bundles.” Here he refers to the fact that theories whose structure groups are certain exceptional Lie groups, such as $E_{6}$ (see Section 5.2), explicitly sever the link to the automorphism groups of $V$.
Thus, methodologically, once you consider the space of more general theories, as Skinner admits, the link between the VB-POV and the PFB-POV is severed; and once it is severed, the door reopens to all the examples of this section—including the slack world itself.
Summing up: upon reconstructing symmetry groups, the VB-POV insists that each gauge subgroup factor be the automorphism group of its corresponding fundamental vector bundle, as represented in Equation (5.15). I do not see this tight identification as a limitation of the VB-POV; on the contrary, it is its chief virtue. By grounding group actions in geometry, the VB-POV rules out many possible theories. That restriction is methodologically valuable, so long as it still encompasses our best physics—and in this case, it does.
@@ 6 Conclusions
Feynman’s Nobel Prize lecture, with which I began, reflected on his alternative formulation of quantum electrodynamics via path integrals. That formulation, like Minkowski’s introduction of spacetime—and indeed many other mathematically equivalent yet conceptually transformative innovations scattered through the history of physics—proved invaluable. They provided new explanations and opened new directions for development. [^34] See \@BBOPcitep\@BAP\@BBN(Hunt, 2025)\@BBCP for more examples of virtuous reformulations of physical and mathematical theories. I do not expect the geometry-first formulation of gauge theory developed here will ascend to comparable heights, nor do I expect it will become orthodoxy, as Feynman’s and Minkowski’s eventually did.
@@
But I do not want to understate what has been gained. The geometry-first formulation postulates a different ontology, offers independent ‘symmetry-free’ explanations of familiar mechanisms in gauge theory, and, above all, eliminates the slack between symmetry and geometry. That tighter fit might be taken to explain why certain group–geometry correspondences—correspondences between properties of vector bosons and matter fermions—are realised in our world and others are not.
Let me summarise the new ‘symmetry-free’ gloss on the explanations involved in the standard Higgs mechanism.
In brief, here is how geometry performs part of the explanatory work of the Higgs mechanism without mentioning symmetry:
(H.i) Since we directly represent the space of non-isomorphic sections of the vector bundle from the start, Goldstone modes never appear here, and so never require elimination (see footnote LABEL:ftnt:_GB).
(H.ii) What in the symmetry-first formulation is described as the acquisition of mass by vector bosons is nothing more than the run-of-the-mill idea that the kinetic term of any section of the vector bundle depends on the affine structure of the vector bundle. The shape operator, or ‘extrinsic curvature’ associated to the direction of the Higgs is the part of the affine structure that ‘gets a mass’. [^35] In the linearised regime, no Killing form is necessary in order to talk about the values of these masses: the inner product on the vector bundles suffices to give dimensions to the linearised perturbations of the flat covariant derivatives.
(H.iii) The covariant derivative along a single section of a vector bundle does not depend on all the affine degrees of freedom of the bundle (for $\text{dim}(E_{x})\geq 2$). Namely, the affine structure that does not appear in the extrinsic curvature remains ‘massless’. The absent degrees of freedom correspond, in the symmetry-first idiom, to the unbroken gauge group, giving rise at the linearised regime to the massless photons.
In this formulation, then, talk of ‘mass acquisition’ may strike a geometry-first militant—say, a relativist—as misplaced. From the VB-POV, what would it even mean for the affine structure, or the covariant derivative $\nabla$, to ‘acquire mass’? Notwithstanding such caution, these assertions acquire a clear meaning in the linearised regime. [^36] To be sure, some would hesitate to say that gravitons acquire mass merely because a spacetime, or a collection thereof, admits a kinetic term for a vector field of constant norm; yet, that they do acquire mass is precisely the consensus for such theories \@BBOPcitep\@BAP\@BBN(Jacobson, 2008)\@BBCP.
Turning to the Yukawa mechanism, discussed in Section 4: I argued that standard presentations are explanatorily ‘opaque,’ and offered instead a more transparent geometric version of the Yukawa form itself. I readily admit that my judgements about what is opaque may stem from a general preference for geometric explanations, simpliciter. But the point remains: as emphasised in Section 2.2, the mere availability of a geometric argument that bypasses representation theory is grist to my mill. The aim, after all, is to open the subject to a different community, with different ways of thinking.
In this spirit, the VB-POV brings out certain features of the Yukawa mechanisms that, to my knowledge, have not been emphasised in the literature. (That does not mean they are controversial; perhaps they are simply too minor to warrant mention in standard presentations.) It also raises new questions.
(Y.i) The left-handed up and down quarks, and likewise the electron and electron-neutrino, are not distinct particles at all: they are simply the components of the (first-generation) left-handed quark fields and leptons lying parallel and orthogonal to the Higgs direction. [^37] That they cannot represent physically distinct particles is noted in some textbooks—cf. \@BBOPcite\@BAP\@BBN(Tong, 2025, p. 185)\@BBCP—but this is usually described in the context of unification ‘before symmetry breaking’. In the the current picture, these just are the parallel/orthogonal components of the same field relative to the Higgs section, and it seems misleading to conflate them with independent particles fields, as they would necessarily swing together under variation.
(Y.ii) In the VB-POV the Yukawa form is geometrically natural. It is the fibrewise contraction to a scalar density built from the intrinsic structures already present on the fundamental bundles.By contrast, in the PFB-POV one postulates a Yukawa map between associated bundles to manufacture invariants while absorbing ambiguities into coupling constants.
@@
(Y.iii) The geometric formulation also sharpens certain physical questions. In particular, the quark Yukawa term depends essentially on the orientation of $\mathbb{C}^{2}$: up-type quarks couple to $\varphi_{c}$, which encodes the oriented area orthogonal to the Higgs field. This explains why the geometry-first picture naturally singles out $SU(2)$ rather than $U(2)$. By contrast, for $\mathbb{C}^{3}$ there appears to be no analogous mechanism: why does the Standard Model employ $SU(3)$ rather than U(3)U(3)? [^38] There is a canonical isomorphism $U(3)\simeq(SU(3)\times U(1))/\mathbb{Z}^{3}$, but the representations of $U(1)$ in the Standard Model do not seem to realise this isomorphism. Benjamin Muntz (p.c.) has suggested that the place to look may be in the triality constraints on baryon coupling: colourless states built from three quarks would not be invariant under the full $U(3)$. Indeed, naively, from the structure endowed to $\mathbb{C}^{3}$ we would expect scalars built of quark–anti-quark pairs, via $\langle\,.\,$$\rangle,$ and others built of three quarks, via $\epsilon$, and nothing more.
Such open questions are emblematic of the broader methodological moral, developed in Section 5. Where the PFB-POV tolerates a loose fit between the structure group and the geometry of its associated bundles, the VB-POV requires each gauge-group factor to coincide with the automorphism group of a fundamental vector bundle (as in Equation (5.15)). In the PFB-POV this looseness gives the gauge group an independent physical role: it can differ from the automorphism group relevant to the matter fields, allowing a certain independence between vector bosons and matter fermions. The VB-POV removes this independence by binding the properties of bosons and fermions to a common geometric source.
But it is important to note that, although the VB-POV tightens the slack between symmetry and geometry, there remains a clear explanatory direction: the group is subservient to the geometry. A striking illustration is provided by the quantisation of charge. In the standard, symmetry-first picture, the fact that charge necessarily comes in discrete values is traced to the topology of a compact group such as $U(1)$. But for non-compact groups, representations could in principle have continuous labels. In the geometry-first picture, by contrast, discretisation follows from the discrete tensorial structure through which matter fields are assembled from the fundamental bundles. Even if the relevant group were non-compact, the tensorial construction would still yield charges labelled by integers.

### v5

@@ 5.3 Recovery of the gauge group from matter bundles
From the PFB-POV this is no defect: the standard strategy assigns each particle type to a section of some associated bundle, without attempting to reconstruct the full gauge group from the geometry of any single fibre. This slack poses no internal difficulty, for the two ingredients—the symmetry and the vector bundle structure—are jointly posited and mutually consistent. [^32] If you think that is not how we actually use the principal-associated bundle formalism, I agree; see Section 5.4 below. But if one tries to recover the group from the automorphisms of those associated bundles, the recovery is not merely incomplete; in general, it is inconsistent.
@@
For concreteness, take $V\simeq{\mathbb{C}}^{3}\otimes{\mathbb{C}}^{2}$, $G=SU(3)\times SU(2)$ and consider two representations: $\rho_{1}=\mathbf{3}\otimes\mathbf{1}$ (a colour triplet, singlet under $SU(2)$), and $\rho_{2}=\mathbf{1}\otimes\mathbf{2}$ (a weak doublet, singlet under $SU(3)$). Thus in this case we would recover $G^{\prime}_{1}=SU(3)$ from one bundle, and $G^{\prime}_{2}=SU(2)$ from the other. What we would like, of course, is $SU(3)\times SU(2)$; but the product structure is nowhere to be found at the level of any single associated bundle.
One might hope to do better by considering the entire matter content at once. Suppose we form the direct sum of all the associated bundles,
$E_{\text{tot}}=E_{1}\oplus E_{2}\oplus\cdots\oplus E_{k},$ (5.12)
and try to recover $G$ from the structure of $E_{\text{tot}}$. There are two natural strategies; both fail.
The first strategy examines the full automorphism group $\mathsf{Aut}(E_{\text{tot}})$. This is too large: it includes not only block-diagonal transformations that act independently on each summand, but also off-diagonal maps that mix summands—transformations with no counterpart in gauge theory.
The second strategy restricts to block-diagonal automorphisms: the subgroup of $\mathsf{Aut}(E_{\text{tot}})$ that preserves each summand $E_{i}$ separately. Call this $\prod_{i}\mathsf{Aut}(E_{i})$. Assuming $G$ acts faithfully on the total matter content (as it does in the Standard Model up to the $\mathbb{Z}_{6}$ kernel), $G$ embeds injectively into $\prod_{i}\mathsf{Aut}(E_{i})$, so no factor of $G$ is invisible. But knowing that $G$ lives inside this product does not tell you which subgroup it is. And crucially, the product decomposition that the block-diagonal strategy hands you is indexed by particle species labeled by ii—one $i$ per associated bundle—whereas the gauge group is a product indexed by interactions: $SU(3)$ for colour, $SU(2)$ for weak isospin, $U(1)$ for hypercharge. These are entirely different decompositions. The species-indexed product $\prod_{i}\mathsf{Aut}(E_{i})$ contains the interaction-indexed product $G=G_{1}\times G_{2}\times G_{3}$ as a subgroup, but does not single it out.
To illustrate: a left-handed quark lives in $E_{Q}=\mathbb{C}^{3}\otimes\mathbb{C}^{2}\otimes\mathbb{C}$, with $\mathsf{Aut}(E_{Q})\simeq U(6)$ (preserving the Hermitian structure on the six-dimensional total fibre). The gauge group $SU(3)\times SU(2)\times U(1)$ embeds in $U(6)$, but so do many other groups. Nothing intrinsic to $E_{Q}$ singles out the tensor-product decomposition $\mathbb{C}^{3}\otimes\mathbb{C}^{2}\otimes\mathbb{C}$ from which the product structure of $G$ derives. That decomposition is additional data—it is the specification of fundamental bundles.
Claim. Given a finite collection of Hermitian vector bundles $\{(E_{i},\langle\cdot,\cdot\rangle_{i})\}_{i=1}^{k}$, or even the enriched collection $\{(E_{i},\langle\cdot,\cdot\rangle_{i},\nabla^{i})\}_{i=1}^{k}$ of bundles equipped with compatible covariant derivatives, the product decomposition of the gauge group $G=\prod_{\alpha}G_{\alpha}$ into independent interaction sectors is not uniquely determined—even when $G$ acts faithfully on the total matter content.
5.4 Fundamental bundles and the Standard Model
The VB-POV avoids the failure we just witnessed, to reconstruct the full product group $G$ from composite bundles, by working at a different level. One does not try to recover the full gauge group from composite bundles. Instead, each factor of $G$ arises as the automorphism group of a fundamental bundle, just as in (2.11). The composite bundles where particles live—tensor products like $E^{3}\otimes E^{2}\otimes E^{1}$—are derived from these. [^33] Of course, particles may have no component in a given fundamental direction. But the gauge group is not recovered from the composite fibres; it is built into the fundamental ones from the start.
As it happens, the symmetry group realised in Nature fits the VB-POV’s stricter requirements. An instructive wrinkle illustrates how.
The nominal gauge group of the Standard Model is $SU(3)\times SU(2)\times U(1)$. A subgroup $\mathbb{Z}_{6}\subset SU(3)\times SU(2)\times U(1)$ acts trivially on the entire particle content of the Standard Model—not merely on individual multiplets, but on every representation that appears. [^34] The $\mathbb{Z}_{6}$ kernel arises from the specific hypercharge assignments of the Standard Model particles. Elements of $SU(3)\times SU(2)\times U(1)$ of the form $(e^{2\pi i/3}\mathds{1}_{3},\,-\mathds{1}_{2},\,e^{i\pi/3})$ and their powers act trivially on all Standard Model fields. The group that acts faithfully on the total matter content is therefore the quotient $SU(3)\times SU(2)\times U(1)/\mathbb{Z}_{6}$.
From the standpoint of the PFB-POV, this is a genuine ambiguity: one can take either the product or the quotient as the structure group, and only global topological data or further physical input can decide between them. From the standpoint of the VB-POV, by contrast, there is nothing to decide. One begins with three fundamental vector bundles—fibres $\mathbb{C}^{3}$, $\mathbb{C}^{2}$, and $\mathbb{C}$, with Hermitian inner products and, for the first two, orientations. Their automorphism groups are $SU(3)$, $SU(2)$, and $U(1)$, and the reconstructed gauge group is their product. The $\mathbb{Z}_{6}$ kernel is a downstream consequence of the specific tensor products nature happens to employ—a curiosity about which multiplets appear in the Standard Model, not a structural problem for the formalism. The analogy is with $SO(3)$ acting on $\mathbb{R}^{3}$: certain tensors built from $\mathbb{R}^{3}$ may be insensitive to certain rotations, but this does not make $SO(3)$ ‘too big’ for the vector space.
If one instead insists that the reconstructed group act faithfully on the total particle content, one obtains $SU(3)\times SU(2)\times U(1)/\mathbb{Z}_{6}$. But this quotient is not a product of classical groups acting on standard spaces in any transparent way, and the geometric interpretation that motivates the frame-bundle picture is lost. A faithfulness requirement, far from cleaning up the recovery, makes it less straightforward.
One might wonder whether grand unified theories fare better. In an $SU(5)$ GUT, for instance, there is a single fundamental bundle with fibre $\mathbb{C}^{5}$, and the entire particle content arises tensorially from it. Here recovery proceeds smoothly—a cleaner fit for the PFB-POV (for the VB-POV the gain is more humble). But $SU(5)$ unification, as far as we know, has failed empirically. The Standard Model, with its product group, its multiple fundamental bundles, and its $\mathbb{Z}_{6}$ kernel, is what we have. The geometry-first picture handles it—and a telling fact about practice suggests that physicists already know this.
Recall the slack world of Section 5.1.a. That is not our world. And barring the standard treatment of the vacuum Yang-Mills case, nor is such a world usually included in the menu of physical possibilities. We do not often seriously consider a mismatch between the symmetries of the gauge sector and those of the matter sector. I claim that there is a tacit assumption built into actual practice. Namely, that each factor of $G$ coincides with the automorphism group of a fundamental bundle:
$G=\prod_{i}G_{i},\quad G_{i}\simeq\rho(G_{i})\simeq\mathsf{Aut}(V_{i}).$ (5.13)
This is precisely what the VB-POV makes explicit.
Now, I admit physicists are not always consistent about this tacit assumption—being tacit makes it harder to see. After describing the simple case where $G\simeq\mathsf{Aut}(V)$, textbooks typically move to composite cases without noting that the product structure must be stipulated. \@BBOPcitet\@BAP\@BBNSkinner (2007)\@BBCP is unusually explicit. On p. 141: “Vector bundles are of relevance to physics because a charged matter field is a section of an associated vector bundle.” After constructing principal bundles from frame bundles—assuming $G\simeq\mathsf{Aut}(V)$—he adds (p. 142): “The most common Lie groups that arise in physics are indeed matrix Lie groups [of that sort], so the two viewpoints are equivalent.” But immediately: “However, in some exotic theories (especially string theory and some grand unified theories) exceptional Lie groups such as $E_{6}$ play an important role, so the fundamental picture is really that of principal bundles.”
The qualification is telling. Skinner restricts equivalence to theories that “commonly arise in physics”—precisely those satisfying the VB-POV’s constraints. What distinguishes those theories? Classical groups acting faithfully in fundamental representations, with matter built tensorially from a small set of fundamental bundles? If those conditions are what make a theory viable, the restriction is not incidental—it is the geometry-first picture in all but name.
@@ 6 Conclusions
6.1 Summary
@@
The geometry-first formulation does three things. It posits a different ontology. It offers independent, symmetry-free explanations of familiar mechanisms. And it eliminates the slack between symmetry and geometry that the principal-bundle formulation tolerates. That tighter fit may help explain why certain group–geometry correspondences are realised in our world and others are not.
What changes in the underlying picture is visible even in the simplest case. In vacuum Yang–Mills theory, the fundamental dynamical object is not a connection on a principal bundle but a covariant derivative on a vector bundle—an affine structure whose automorphism group happens to be a Lie group, but whose description need never mention one.
Consider how geometry does the explanatory work of the Higgs mechanism without mentioning symmetry. Since we represent non-isomorphic sections directly from the start, Goldstone modes never appear and never require elimination (see footnote 18). What the symmetry-first formulation calls ‘mass acquisition’ is simply this: the kinetic term of any section depends on the bundle’s affine structure, and the shape operator along the Higgs direction—the extrinsic curvature—is the part that acquires mass. [^35] In the linearised regime, no Killing form is needed: the inner product on vector bundles suffices to give dimensions to perturbations of flat covariant derivatives. But the covariant derivative along a single section does not exhaust the affine degrees of freedom (when $\text{dim}(E_{x})\geq 2$). The degrees absent from the extrinsic curvature remain massless—corresponding, in symmetry-first language, to the unbroken gauge group and its massless photons. Strictly speaking, ‘mass acquisition’ is a misnomer from the geometry-first standpoint; what could it mean for a covariant derivative to acquire mass? But the phrase acquires clear meaning in the linearised regime. [^36] Some would hesitate to say gravitons acquire mass merely because a spacetime admits a kinetic term for a constant-norm vector field; yet that they do is the consensus for such theories (\@BBOPcitep\@BAP\@BBN(Jacobson, 2008)\@BBCP).
@@
The Yukawa mechanism, too, gains clarity. The Yukawa form becomes geometrically natural—a fibrewise contraction built from structures already present on the fundamental bundles—rather than a representation-theoretic construction that postulates maps between associated bundles and absorbs ambiguities into coupling constants. And the formulation sharpens physical questions: the quark Yukawa term depends essentially on the orientation of $\mathbb{C}^{2}$, which explains why $SU(2)$ rather than $U(2)$ appears naturally. For $\mathbb{C}^{3}$, by contrast, I don’t know the analogous mechanism—why does the Standard Model employ $SU(3)$ rather than U(3)U(3)? [^37] There is a canonical isomorphism $U(3)\simeq(SU(3)\times U(1))/\mathbb{Z}^{3}$, but the $U(1)$ representations in the Standard Model do not realise it. Benjamin Muntz (p.c.) suggests looking at triality constraints on baryon coupling: colourless three-quark states would not be invariant under $U(3)$.
The group is subservient to the geometry, and this changes familiar explanations. Charge quantisation (Section 5.2) is a striking example: discreteness follows from the tensorial construction of matter fields, not from the topology of a compact group. Even if the automorphism group of the fibre were non-compact, the charge lattice would persist—its generator fixed by the choice of fundamental bundle, not by a topological identification in the group.
6.2 Two defences